In Kendrick's dream home, he is now putting a sidewalk that goes from his driveway to his front porch. He is pouring 4 inches deep. The height is 10 ft, and the length is 12 ft. The concrete will cost him $85 per cubic yard. How much will it cost to pour the sidewalk? Thank you for helping me.
Answer by checkley79(3341)

4/12*10*12=1/3*120=40 cubic feet of concrete.
3*3*3=27 cubic feet per yard
40/27=1.48 cubic yards.
1.48*85=$125.80 cost to pour the concrete.